# From Dormant to Deleted: Tamper-Resistant Unlearning Through Weight-Space Regularization

### v1

@@ 1 Introduction
Machine unlearning is the problem of removing the influence of specific training datapoints, the forget set, from a pretrained model. This was initially motivated from the perspective of privacy, the right-to-be-forgotten (cao2015towards, ; neel2021descent, ) and data protection policies (mantelero2013eu, ), and recently applied to a range of problems, including removing harmful knowledge (li2024wmdp, ; zou2024circuitbreakers, ; eldan2023whosharry, ).
@@
In this work, we take a different perspective, focusing on relearning attacks on unlearned models. We consider tampering attacks (che2025tamperingattacks, ; tamirisa2408tamper, ), where the attacker is able to fine-tune the model weights. This direction is inspired by a growing set of observations in the context of unlearning “knowledge” in LLMs: fine-tuning an unlearned model even on seemingly benign data may cause the believed-to-be-unlearned knowledge to re-emerge (hu2024jogging, ; lucki2024adversarial, ; deeb2024unlearning, ; che2025tamperingattacks, ). However, these studies are carried out under conditions that make it difficult to draw clear conclusions. First, these works study unlearning knowledge, capabilities, or topics, where the problem is inherently under specified. For example, given a dataset containing knowledge necessary for making bioweapons, the goal may be to fully remove the capability of constructing bioweapons, while preserving general knowledge of biology. In this setting, it is hard to draw a clear line between forbidden and permissible knowledge and pinpoint all training examples responsible for acquiring different types of knowledge. Furthermore, knowledge is hard to measure, especially due to nuances of natural language, requiring question-answer evaluation which may be sensitive to the particular phrasing. Finally, because LLMs can make complex inferences beyond the training set, not all knowledge that is extractable should be attributed to a failure of unlearning shumailov2024ununlearning . In many cases, the acquisition of knowledge is natural and unavoidable, hence, making it difficult to distinguish between the two in these problem settings.
@@
To address these issues, we study relearning attacks in a setting that allows for controlled experimentation: unlearning specific training examples from (small) vision classification models, a problem for which a plethora of approximate unlearning methods have been developed and tested (golatkar2021mixed, ; graves2021amnesiac, ; goel2022towards, ; kurmanji2024scrub, ; foster2024ssd, ; torkzadehmahani2024improved, ; zhao2024makes, ; sepahvand2025selective, ). In example-level unlearning, the gold standard is clear, namely, to retrain the model without the forget set. As our models are small enough, we can compute the gold-standard solution and use it as the reference point for comparison. Because we consider classification, we can also use accuracy as a simple and well-understood measure of performance. These properties combined allow us to compare the tamper resistance of different unlearning algorithms with the correct reference point, and therefore draw conclusions about their quality.
@@
We evaluate a range of increasingly-complex unlearning algorithms in this setting and discover a surprising finding: for numerous unlearning algorithms, the accuracy of the forget set jumps from around 50% post-unlearning to nearly 100% after fine-tuning the unlearned models on only the retain set, which is disjoint from the forget set. Fig. 1 shows this phenomenon on CIFAR-10 using ResNet-18, after having attempted to unlearn a subset of atypical instances of class ‘airplane’. We note that a model retrained from scratch without the forget set does not exhibit this behaviour, with the accuracy remaining at 50%. Therefore, the recovery of forget set accuracy can be safely interpreted as a failure of these algorithms to fully remove the influence of the data points in the forget set.
@@ 2 Background and Related Work
Unlearning. The problem of machine unlearning was introduced by (cao2015towards, ). The goal is to remove the influence of a “forget set” from a model that was trained on a dataset including that set. This was motivated by privacy and right-to-be-forgotten policies (mantelero2013eu, ). The perfect unlearning method, from the perspective of fully erasing the influence of the forget set, is to simply retrain the model excluding that set. However, the computationally prohibitive training costs make such an approach infeasible in most practical cases. bourtoule2021shardedtrain propose to shard the dataset, and train an ensemble model over it, allowing to selectively retrain only the affected parameters. However, the computational cost is still prohibitive in the worst case, while also leading to poorer performance in some cases due to the use of specialized architectures. These issues motivated the development of approximate methods that accept imperfect unlearning in exchange for greater efficiency. This is the family of methods we focus on in this work. The goal in approximate unlearning is to post-process the trained model as efficiently as possible in order to closely match the model which is retrained from scratch using only a small amount of model fine-tuning (jia2023l1sparse, ; chundawat2023badteacher, ; foster2024ssd, ; kurmanji2024scrub, ; li2024wmdp, ; zou2024circuitbreakers, ; zhao2024unlearninghard, ; triantafillou2024unlearningcomp, ). This is a challenging problem as imperfect attempts to erase the influence of the forget set post-hoc may have a number of unwanted side-effects, such as harming the overall utility of the model (triantafillou2024unlearningcomp, ).
Unlearning quality metrics. Since most approximate unlearning methods that are applicable to deep models do not come with theoretical guarantees about the quality of their approximation, we are required to estimate how well they approximate retraining from scratch empirically. This is a research problem in and of itself, and current rigorous metrics are very computationally expensive triantafillou2024unlearningcomp ; hayes2024inexact . Furthermore, unlearning entails fundamental trade-offs, such as between forgetting and maintaining the model’s utility. This requires a multifaceted evaluation metric that captures relevant factors aside from forgetting quality. Commonly, in vision classification, the accuracy on the retain set and the accuracy on the test set are used to measure model utility. In similar spirit to our work, time to relearn (golatkar2021mixed, ) quantifies the strength of unlearning by the number of optimization steps required to reacquire forgotten information by directly fine-tuning on it. We instead show that we can restore forget set accuracy even if fine-tuning only on a subset of it, or even only the retain set.
@@ 4 Experimental Setup
Pretraining We pretrain the model for 300 epochs using Adam optimizer (kingma2014adam, ) with a learning rate of 1e-4, cosine learning rate decay with a decay factor of 0.1, batch size of 128, and a weight decay of 1e-4 in all configurations.
@@
Relearning attack. During this phase, we fine-tune on a combination of the retain set $\mathcal{D}_{R}$ and a subset of the forget set for relearning ($\mathcal{D}_{F_{re}}$). We explore the impact of different choices for relearning examples in Appendix F. We again use a small learning rate of 1e-5 without any weight decay, and optimize the model for just 10 epochs (except Fig. 1 where we optimized the model for 300 epochs). Similar to the pretraining stage, we use a cosine learning rate decay with a decay factor of 0.1.
@@ 4.1 Baseline Unlearning Methods
Figure 2: Scatter-plots with test-set accuracy on the abscissa and accuracy on the held-out portion of the forget set, $\mathcal{D}_{F_{ho}}$, on the ordinate. The left-most subplot indicates performance immediately following unlearning. The next three subplots are following a relearning attack with instances of the retain set, $\mathcal{D}_{R}$ and a varying number of instances of the forget set, $\mathcal{D}_{F_{re}}$ (0, 10, and 100, respectively). Each point is the average performance of the last 50 steps (see Fig. 10 for the whole trajectory for sub-class unlearning and Fig. 11 for the trajectory for class-agnostic unlearning). The forget set is comprised of atypical examples (from the ‘airplane’ class i.e., sub-class unlearning for the top row and all classes i.e., class-agnostic unlearning in the bottom row) in CIFAR-10. The figure indicates that many methods achieve near-perfect recovery of unlearned knowledge with only a small amount of model fine-tuning, even with 0 relearning examples (fine-tuning on only the retain set). Weight Distortion, CBFT, and Weight Dist Reg are introduced in Section 6.
@@
Random Relabeling graves2021amnesiac relabels every example in the forget set with a random label. Note that we used an aggressive version of random relabeling where we re-assign a new label at every fine-tuning step. Hence, this can be considered analogous to minimizing divergence to a uniform distribution.
@@ 5 Recent Unlearning Algorithms are not Tamper-Resistant
The role of typicality for learning and unlearning. Not all examples are equally easy to unlearn, and not all examples are equally easy to (re)learn. Typical examples can be particularly easy to unlearn as even a no-op unlearning already yields similar predictions on these examples as the retrained from scratch model (jia2023l1sparse, ). This is because they are, by definition, easy to predict regardless of whether or not they are part of the training set  (feldman2020neural, ; jiang2020characterizing, ; baldock2021deep, ; siddiqui2022metadata, ). Generally, because typical examples are more likely to be predicted correctly, they are easier to relearn as well. Based on this, we hypothesize that the typicality of examples is a key property that will determine relearning patterns and differences between unlearning algorithms compared to retraining from scratch.
@@
When the forget set consists of typical examples, relearning is trivial and uninteresting. The reason is that accuracy of the retrain-from-scratch model is essentially perfect on the forget set, even before relearning is applied, and all unlearning methods exhibit similar behavior (for the sake of completeness, we show the results for this case in Fig. 6Appendix A). Similarly, when the forget set consists of randomly selected examples, performance of retrain-from-scratch is nearly perfect because randomly selected examples are predominantly typical (feldman2020neural, ; feldman2020does, ; baldock2021deep, ; siddiqui2022metadata, ) (again, for completeness, this case is shown in Fig. 7Appendix B). Consequently, we focus on atypical forget set items for the remainder of the paper; for these items, retrain-from-scratch will not predict correctly prior to relearning and thus we can measure the effectiveness of a relearning attack.
@@
Relearning attacks succeed against several unlearning baselines. We present the results for sub-class unlearning for the forget set of atypical examples in Fig. 2. In this and subsequent figures, existing methods are indicated by pastel colors. New methods, to be introduced shortly, are represented by saturated colors. For the most part, the existing methods all behave similarly and the reader need not attend to the individual methods. The accuracy of retrain from scratch is less than 50% on that forget set, and remains almost exactly the same when subjected to the relearning attack of fine-tuning only on the retain set (i.e., 0 relearning examples). The accuracy of this model shifts up slightly as we increase the number of relearning examples from the forget set (going from left to right). In stark contrast, the different unlearning methods we evaluate show a qualitatively different trend. We make the striking observation that some methods (such as Circuit Breakers, SCRUB, and Random Relabeling) are very susceptible to relearning attacks. For these methods, forget-set accuracy drops down after unlearning, near the desired reference point of the retrained model. However, upon relearning even on just the retain set, the model achieves near-perfect accuracy on the forget set—a jump from near 50% post-unlearning to nearly 100% after relearning.
@@
Sub-class vs class-agnostic unlearning. We further attempt to understand if such differences exist in a class-agnostic forget-set setting (with the same number of forget-set examples, i.e., 500), which is now comprised of atypical examples across classes. Fig. 2 (bottom) shows that class-agnostic unlearning better differentiates among methods compared to sub-class unlearning. We see a wider spread among methods, even in the first subplot (post-unlearning, pre-relearning), since class-agnostic unlearning is harder. We also observe that while for sub-class unlearning, the performance of retrain-from-scratch (black star) on $\mathcal{D}_{F_{ho}}$ shifts upwards as more relearning examples are used, this shift does not occur with class-agnostic unlearning. This result is expected since relearning on a larger number of atypical examples does not improve performance on a disjoint set of other atypical examples (by definition of atypicality (feldman2020neural, )). There is again a stark contrast between the performance of retrain-from-scratch model on $\mathcal{D}_{F_{ho}}$, where post-relearning accuracy remains very low, and many unlearning algorithms, where post-relearning accuracy again reaches nearly 100%. Importantly, the relative ranking between methods in the sub-class and the class-agnostic case remains consistent.
@@
Relearning attacks succeed even against methods designed for tamper-resistance. We further compare two-phase methods that assume an initial unlearning round, followed by a subsequent training round in order to reduce susceptibility against relearning attacks. This is inspired by the methodology of TAR (tamirisa2408tamper, ), which is explicitly designed to increase the resistance of the unlearned model against fine-tuning-based relearning attacks. Despite TAR’s success in the case of language models, Fig. 3 shows that it struggles to provide any resistance against relearning attacks in our case.
@@ 6 A Weight-Space View on Understanding and Improving Tamper-Resistance
In the previous section, we demonstrated the susceptibility of existing prominent unlearning methods to tampering (Fig. 2). However, it is unclear what makes a method vulnerable or robust to these relearning attacks. In this section, we shed light on this question through a weight-space view. Specifically, we hypothesize that the susceptibility of an unlearned model to relearning may be associated with failing to move ‘far enough’ from the pretrained model in weight-space. We explore this hypothesis from two perspectives: (i) by measuring distances in weight-space, and (ii) through Linear Mode Connectivity analysis (frankle2020linear, ). We then use these tools to interpret the tamper-resistance profiles of previously-proposed unlearning algorithms (Fig. 2) that show better tamper-resistance (such as Catastrophic Forgetting and L1 Sparse) compared to the ones that exhibit worse tamper-resistance (such as Random Relabeling, Circuit Breakers, and SCRUB).
Figure 5: $L_{2}$ norm of the difference between the parameters of the pretrained and the unlearned models induced by different methods. We consider only the difference in the parameters, while ignoring the batch-norm statistics for ResNet-18 trained on CIFAR-10, where the forget set is comprised of atypical examples.
@@
Fig. 5 plots $L_{2}$ distance between the pretrained model and the unlearned model. We see that existing methods that we previously showed to be susceptible to relearning (pastel colors) induce only small movement in parameter space, indicated by the small $L_{2}$ distance.
Generally, we see that methods with higher $L_{2}$ distance exhibit higher robustness against relearning attacks presented in Fig. 2. Notably, two methods based on these insights which we describe shortly, Weight Distortion and Weight Dist Reg, have the highest $L_{2}$-norm and higher tamper-resistance. We remark that, out of previous methods, those with increased tamper-resistance (such as Catastrophic Forgetting and L1 Sparse) have comparatively higher distance norm compared to methods with very poor tamper resistance (such as Random Relabeling, Circuit Breakers, and SCRUB). We note that both Catastrophic Forgetting and L1 Sparse use weight-space regularizers ($L_{2}$ and $L_{1}$, respectively). Fine-tuning without any regularizer failed to unlearn the forget set in our evaluations, which aligns with our weight-space interpretation.
@@
We further investigate the relationship in weight-space between the unlearned and the retrained from scratch model via the lens of LMC (frankle2020linear, ). We plot the accuracy when interpolating between the two models (where we interpolate both model parameters as well as model batch-norm statistics) along a linear path. We compare the pretrained and retrained from scratch model (which are trained from the same initialization), as well as the pretrained and different unlearned models in Fig. 4. Looking at the retrain from scratch plot, we see that there is a high-loss barrier between the two models, meaning that they are not in linearly connected modes. The same holds for methods with some tamper resistance (like Catastrophic Forgetting and L1 Sparse) but not for those vulnerable to tampering attacks like SCRUB, where we observe no such barrier.

### v2

@@ 1 Introduction
Machine unlearning is the problem of removing the influence of specific training datapoints, the forget set, from a pretrained model. This was initially motivated from the perspective of privacy, the right-to-be-forgotten (cao2015towards, ; neel2021descent, ), data protection policies (mantelero2013eu, ), and recently applied to a range of problems, including removing harmful knowledge (li2024wmdp, ; zou2024circuitbreakers, ; eldan2023whosharry, ).
@@
In this work, we take a different perspective, focusing on relearning attacks on unlearned models. We consider tampering attacks (che2025tamperingattacks, ; tamirisa2408tamper, ), where the attacker is able to fine-tune the model weights. This direction is inspired by a growing set of observations in the context of unlearning “knowledge” in LLMs: fine-tuning an unlearned model even on seemingly benign data may cause the believed-to-be-unlearned knowledge to re-emerge (hu2024jogging, ; lucki2024adversarial, ; deeb2024unlearning, ; che2025tamperingattacks, ). However, these studies are carried out under conditions that make it difficult to draw clear conclusions. First, these works study unlearning knowledge, capabilities, or topics, where the problem is inherently under specified. For example, given a dataset containing knowledge necessary for making bioweapons, the goal may be to fully remove the capability of constructing bioweapons, while preserving general knowledge of biology. In this setting, it is hard to draw a clear line between forbidden and permissible knowledge and pinpoint all training examples responsible for acquiring different types of knowledge. Furthermore, knowledge is hard to measure, especially due to nuances of natural language, requiring question-answer evaluation which may be sensitive to the particular phrasing. Finally, because LLMs can make complex inferences beyond the training set, not all knowledge that is extractable should be attributed to a failure of unlearning (shumailov2024ununlearning, ). In many cases, the acquisition of knowledge is natural and unavoidable, hence, making it difficult to distinguish between the two in these problem settings.
@@
To address these issues, we study relearning attacks in a setting that allows for controlled experimentation: unlearning specific training examples from (small) vision classification models, a problem for which a plethora of approximate unlearning methods have been developed and tested (golatkar2021mixed, ; graves2021amnesiac, ; goel2022towards, ; kurmanji2024scrub, ; foster2024ssd, ; torkzadehmahani2024improved, ; zhao2024what, ; sepahvand2025selective, ). In example-level unlearning, the gold standard is clear, namely, to retrain the model without the forget set. As our models are small enough, we can compute the gold-standard solution and use it as the reference point for comparison. Because we consider classification, we can also use accuracy as a simple and well-understood measure of performance. These properties combined allow us to compare the tamper resistance of different unlearning algorithms with the correct reference point, and therefore draw conclusions about their quality.
@@
We evaluate a range of increasingly-complex unlearning algorithms in this setting and discover a surprising finding: for numerous unlearning algorithms, the accuracy of the forget set jumps from around 50% post-unlearning to nearly 100% after fine-tuning the unlearned models on only the retain set, which is disjoint from the forget set. Fig.˜1 shows this phenomenon on CIFAR-10 using ResNet-18, after having attempted to unlearn a subset of atypical instances of class ‘airplane’. We note that a model retrained from scratch without the forget set does not exhibit this behaviour, with the accuracy remaining at 50%. Therefore, the recovery of forget set accuracy can be safely interpreted as a failure of these algorithms to fully remove the influence of the datapoints in the forget set.
@@ 2 Background and Related Work
Unlearning. The problem of machine unlearning was introduced by (cao2015towards, ). The goal is to remove the influence of a “forget set” from a model that was trained on a dataset including that set. This was motivated by privacy and right-to-be-forgotten policies (mantelero2013eu, ). The perfect unlearning method, from the perspective of fully erasing the influence of the forget set, is to simply retrain the model excluding that set. However, the computationally prohibitive training costs make such an approach infeasible in most practical cases. bourtoule2021shardedtrain propose to shard the dataset, and train an ensemble model over it, allowing to selectively retrain only the affected parameters. However, the computational cost is still prohibitive in the worst case, while also leading to poorer performance in some cases due to the use of specialized architectures. These issues motivated the development of approximate methods that accept imperfect unlearning in exchange for greater efficiency. This is the family of methods we focus on in this work. The goal in approximate unlearning is to post-process the trained model as efficiently as possible in order to closely match the model which is retrained from scratch using only a small amount of model fine-tuning (jia2023l1sparse, ; chundawat2023badteacher, ; foster2024ssd, ; kurmanji2024scrub, ; li2024wmdp, ; zou2024circuitbreakers, ; zhao2024what, ; triantafillou2024unlearningcomp, ). This is a challenging problem as imperfect attempts to erase the influence of the forget set post-hoc may have a number of unwanted side-effects, such as harming the overall utility of the model (triantafillou2024unlearningcomp, ).
Unlearning quality metrics. Since most approximate unlearning methods that are applicable to deep models do not come with theoretical guarantees about the quality of their approximation, we are required to estimate how well they approximate retraining from scratch empirically. This is a research problem in and of itself, and current rigorous metrics are very computationally expensive (triantafillou2024unlearningcomp, ; hayes2024inexact, ). Furthermore, unlearning entails fundamental trade-offs, such as between forgetting and maintaining the model’s utility. This requires a multifaceted evaluation metric that captures relevant factors aside from forgetting quality. Commonly, in vision classification, the accuracy on the retain set and the accuracy on the test set are used to measure model utility. In similar spirit to our work, time to relearn (golatkar2021mixed, ) quantifies the strength of unlearning by the number of optimization steps required to reacquire forgotten information by directly fine-tuning on it. We instead show that we can restore forget set accuracy even when fine-tuning on only a subset of it, or solely on the retain set.
@@ 4 Experimental Setup
Pretraining. We pretrain the model for 300 epochs using Adam optimizer (kingma2014adam, ) with a learning rate of 1e-4, cosine learning rate decay with a decay factor of 0.1, batch size of 128, and a weight decay of 1e-4 in all configurations.
@@
Relearning attack. During this phase, we fine-tune on a combination of the retain set $\mathcal{D}_{R}$ and a subset of the forget set for relearning ($\mathcal{D}_{F_{re}}$). We explore the impact of different choices for relearning examples in Appendix˜G. We again use a small learning rate of 1e-5 without any weight decay, and optimize the model for just 10 epochs (except Fig.˜1 where we optimized the model for 300 epochs). Similar to the pretraining stage, we use a cosine learning rate decay with a decay factor of 0.1.
@@ 4.1 Baseline Unlearning Methods
Figure 2: Scatter-plots with test-set accuracy on the x-axis and accuracy on the held-out portion of the forget set, $\mathcal{D}_{F_{ho}}$, on the y-axis. The left-most subplot indicates performance immediately following unlearning. The next three subplots are following a relearning attack with instances of the retain set, $\mathcal{D}_{R}$ and a varying number of instances of the forget set, $\mathcal{D}_{F_{re}}$ (0, 10, and 100, respectively). Each point is the average performance of the last 50 steps (see Fig.˜11 for the whole trajectory for sub-class unlearning and Fig.˜12 for the trajectory for class-agnostic unlearning). The forget set is comprised of atypical examples (from the ‘airplane’ class, i.e., sub-class unlearning for the top row and all classes, i.e., class-agnostic unlearning in the bottom row) in CIFAR-10. The figure indicates that many methods achieve near-perfect recovery of unlearned knowledge with only a small amount of model fine-tuning, even with 0 relearning examples (fine-tuning on only the retain set). Weight Distortion, CBFT, and Weight Dist Reg are introduced in Section˜6.
@@
Random Relabeling (graves2021amnesiac, ) relabels every example in the forget set with a random label. Note that we used an aggressive version of random relabeling where we re-assign a new label at every fine-tuning step. Hence, this can be considered analogous to minimizing divergence to a uniform distribution.
@@ 5 Recent Unlearning Algorithms are not Tamper-Resistant
The role of typicality for learning and unlearning. Not all examples are equally easy to unlearn, and not all examples are equally easy to (re)learn. Typical examples can be particularly easy to unlearn as even a no-op unlearning already yields similar predictions on these examples as the retrained from scratch model (jia2023l1sparse, ). This is because they are, by definition, easy to predict regardless of whether they are part of the training set or not (feldman2020neural, ; jiang2020characterizing, ; baldock2021deep, ; siddiqui2022metadata, ). Generally, because typical examples are more likely to be predicted correctly, they are easier to relearn as well. Based on this, we hypothesize that the typicality of examples is a key property that will determine relearning patterns and differences between unlearning algorithms compared to retraining from scratch.
@@
When the forget set consists of typical examples, relearning is trivial and uninteresting. The reason is that accuracy of the retrain-from-scratch model is essentially perfect on the forget set, even before relearning is applied, and all unlearning methods exhibit similar behavior (for the sake of completeness, we show the results for this case in Fig.˜7Appendix˜A). Similarly, when the forget set consists of randomly selected examples, performance of retrain-from-scratch is nearly perfect because randomly selected examples are predominantly typical (feldman2020neural, ; feldman2020does, ; baldock2021deep, ; siddiqui2022metadata, ) (again, for completeness, this case is shown in Fig.˜8Appendix˜B). Consequently, we focus on atypical forget set items for the remainder of the paper; in this case, retrain-from-scratch will not predict correctly prior to relearning and thus we can measure the effectiveness of a relearning attack.
@@
Relearning attacks succeed against several unlearning baselines. We present the results for sub-class unlearning for the forget set of atypical examples in Fig.˜2 (top). In this and subsequent figures, existing methods are indicated by pastel colors. New methods, to be introduced shortly, are represented by saturated colors. For the most part, the existing methods all behave similarly and the reader need not attend to the individual methods. The accuracy of retrain from scratch is less than 50% on that forget set, and remains almost exactly the same when subjected to the relearning attack of fine-tuning only on the retain set (i.e., 0 relearning examples). The accuracy of this model shifts up slightly as we increase the number of relearning examples from the forget set (going from left to right). In stark contrast, the different unlearning methods we evaluate show a qualitatively different trend. We make the striking observation that some methods (such as Circuit Breakers, SCRUB, and Random Relabeling) are very susceptible to relearning attacks. For these methods, forget-set accuracy drops down after unlearning, near the desired reference point of the retrained model. However, upon relearning even on just the retain set, the model achieves near-perfect accuracy on the forget set—a jump from near 50% post-unlearning to nearly 100% after relearning.
@@
Sub-class vs class-agnostic unlearning. We further attempt to understand if such differences exist in a class-agnostic forget-set setting (with the same number of forget-set examples, i.e., 500), which is now comprised of atypical examples across classes. Fig.˜2 (bottom) shows that class-agnostic unlearning better differentiates among methods compared to sub-class unlearning. We see a wider spread among methods, even in the first subplot (post-unlearning, pre-relearning), since class-agnostic unlearning is harder. We also observe that while for sub-class unlearning, the performance of retrain-from-scratch (black star) on $\mathcal{D}_{F_{ho}}$ shifts upwards as more relearning examples are used, this shift does not occur with class-agnostic unlearning. This result is expected since relearning on a larger number of atypical examples does not improve performance on a disjoint set of other atypical examples (by definition of atypicality (feldman2020neural, )). There is again a stark contrast between the performance of retrain-from-scratch model on $\mathcal{D}_{F_{ho}}$, where post-relearning accuracy remains very low, and many unlearning algorithms, where post-relearning accuracy again reaches nearly 100%. Importantly, the relative ranking between methods in the sub-class and the class-agnostic case remains consistent.
@@
Relearning attacks succeed even against methods designed for tamper-resistance. We further compare two-phase methods that assume an initial unlearning round, followed by a subsequent training round in order to reduce susceptibility to relearning attacks. This is inspired by the methodology of TAR (tamirisa2408tamper, ), which is explicitly designed to increase the resistance of the unlearned model against fine-tuning-based relearning attacks. Despite TAR’s success in the case of language models, Fig.˜3 shows that it struggles to provide any resistance against relearning attacks in our case.
@@ 6 A Weight-Space View on Understanding and Improving Tamper-Resistance
In the previous section, we demonstrated the susceptibility of existing prominent unlearning methods to tampering (Fig.˜2). However, it is unclear what makes a method vulnerable or robust to these relearning attacks. In this section, we shed light on this question through a weight-space view. Specifically, we hypothesize that the susceptibility of an unlearned model to relearning may be associated with failing to move ‘far enough’ from the pretrained model in weight-space. We explore this hypothesis from two perspectives: (i) by measuring distances in weight-space, and (ii) through Linear Mode Connectivity analysis (frankle2020linear, ). We then use these tools to interpret the tamper-resistance profiles of previously-proposed unlearning algorithms (Fig.˜2) that show better tamper-resistance (such as Catastrophic Forgetting and L1 Sparse) compared to the ones that exhibit worse tamper-resistance (such as Random Relabeling, Circuit Breakers, and SCRUB).
Figure 5: $L_{2}$ norm of the difference between the parameters of the pretrained and the unlearned models induced by different unlearning methods. We consider only the difference in the parameters, while ignoring the batch-norm statistics for ResNet-18 trained on CIFAR-10, where the forget set is comprised of atypical examples.
@@
Fig.˜5 plots $L_{2}$ distance between the pretrained model and the unlearned model. We see that existing methods that we previously showed to be susceptible to relearning (pastel colors) induce only small movement in parameter space, indicated by the small $L_{2}$ distance.
Generally, we see that methods with higher $L_{2}$ distance exhibit higher robustness against relearning attacks presented in Fig.˜2. Notably, two methods based on these insights which we describe shortly, Weight Distortion and Weight Dist Reg, have the highest $L_{2}$-norm and higher tamper-resistance. We remark that, out of previous methods, those with increased tamper-resistance (such as Catastrophic Forgetting and L1 Sparse) have comparatively higher distance norm compared to methods with very poor tamper resistance (such as Random Relabeling, Circuit Breakers, and SCRUB). We note that both Catastrophic Forgetting and L1 Sparse use weight-space regularizers ($L_{2}$ and $L_{1}$, respectively). Fine-tuning without any regularizer failed to unlearn the forget set in our evaluations, which aligns with our weight-space interpretation. We consider more elaborate analysis regarding layers that are particularly important for robustness against relearning attacks as an interesting direction for future work.
@@
We further investigate the relationship in weight-space between the unlearned and the retrained from scratch model via the lens of LMC (frankle2020linear, ). We plot the accuracy when interpolating between the two models (where we interpolate both model parameters as well as model batch-norm statistics) along a linear path. We compare the pretrained and retrained from scratch model (which are trained from the same initialization), as well as the pretrained and different unlearned models in Fig.˜4. Looking at the retrain from scratch plot, we see that there is a high-loss barrier between the two models, meaning that they are not in linearly connected modes. The same holds for methods with some tamper resistance (like Catastrophic Forgetting and L1 Sparse) but not for those vulnerable to tampering attacks like SCRUB, where we observe no such barrier.